Nitrogen and hydrogen combine at a high temperature, in the presence of a catalyst, to produce ammonia. N2(g)+3H2(g)⟶2NH3(g) Assume 0.180 mol N2 and 0.564 mol H2 are present initially. How many moles of H2 remain?
Answer #1

Balanced chemical equation is:

N2 + 3 H2 ---> 2 NH3 + CuS

1 mol of N2 reacts with 3 mol of H2

for 0.18 mol of N2, 0.54 mol of H2 is required

But we have 0.564 mol of H2

so, N2 is limiting reagent

According to balanced equation

mol of H2 reacted = (3/1)* moles of N2

= (3/1)*0.18

= 0.54 mol

mol of H2 remaining = mol initially present - mol reacted

mol of H2 remaining = 0.564 - 0.54

mol of H2 remaining = 2.4*10^-2 mol

Answer: 0.0240 mol